OpenStudy (anonymous):

determine whether the series converges absolutely, converges conditionally or diverges

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-1)^n *(2n)! }{ 5n!}\]

OpenStudy (anonymous):

@perl need help with this problem

OpenStudy (anonymous):

the answer is diverge but how do i proof or show it diverges?

OpenStudy (freckles):

try ratio test

OpenStudy (freckles):

by the way is the bottom 5*n! or (5n)!

OpenStudy (freckles):

it doesn't matter we would still use ratio test

OpenStudy (freckles):

but anyways let me know if you need anymore help

OpenStudy (freckles):

useful hints: (n+1)!=(n+1)*n! or you know like (2n+1)!=(2n+1)*(2n)!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

ratio if it is what you wrote, the terms don't go to zero, do they?

OpenStudy (anonymous):

@freckles yes its 5*n!

OpenStudy (anonymous):

@freckles sorry i was away from keyboard

OpenStudy (freckles):

its k still here

OpenStudy (anonymous):

so i have the multiply the numerator and denominator

OpenStudy (freckles):

\[a_n=\frac{(-1)^n(2n)!}{5 \cdot n!} \\ a_{n+1}=\frac{(-1)^{n+1}(2[n+1])!}{5 (n+1)!} =\frac{(-1)^n(-1)(2n+2)!}{5(n+1)!}\\ =\frac{(-1)(-1)^n(2n+2) \cdot (2n+1) \cdot (2n)! }{5(n+1) \cdot n!}\]

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}|\frac{a_{n+1}}{a_n}|=?\]

OpenStudy (anonymous):

\[\frac{ (-1)^n (-1)(2n+2) }{ 5(n+1)! }/\frac{ (-1)^n * (2n)! }{ 5*n! } = \frac{ 10n+10 * n! }{ 10n(n+1)! }\]

OpenStudy (anonymous):

sort of right?

OpenStudy (freckles):

I'm not quite sure how you got that

OpenStudy (anonymous):

its indeterminate so we got to use l'hopital's rule 10+10n!/10n^2+10n!

OpenStudy (freckles):

\[\frac{(-1)\cancel{(-1)^n}(2n+2)(2n+1) \cancel{(2n)!}}{\cancel{5}(n+1) \cancel{n!}} \cdot \frac{\cancel{5} \cancel{n!}}{\cancel{(-1)^n} \cancel{(2n)!}}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

why 2n+1?

OpenStudy (anonymous):

i mean where did you get it from?

OpenStudy (freckles):

remember n!=n(n-1)(n-2)(n-3)! so if you have (2n)!=(2n)*(2n-1)*(2n-2)*(2n-3)*(2n-4)! or if you don't won't to write that many factors (2n)!=(2n)*(2n-1)! so if you have (2n+2)!=(2n+2)*(2n+1)*(2n)*(2n-1)*(2n-2)! or as I wrote it above (2n+2)!=(2n+2)*(2n+1)*(2n)!

OpenStudy (freckles):

like you know 15!=15*14*13*12! 16!=16*15! Like 2n+2=16 when n=7 so 16!=(2*7+2)!=(2*7+2)*(2*7+1)*(2*7)! you are just taking one away from the previous and tacking on a ! if you don't keep writing

OpenStudy (anonymous):

oh okay so i got infinity as the result does that mean we have to use l'hopital's rule?

OpenStudy (freckles):

which part gave you infinity

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} (2n+2)(2n+1)\]

OpenStudy (freckles):

and that is over (n+1) right

OpenStudy (anonymous):

oh yes

OpenStudy (freckles):

but its like you basically have 4n^2/n=4n and yes 4n goes to infinity as n goes to infinity

OpenStudy (anonymous):

and thats diverge

OpenStudy (freckles):

yep

OpenStudy (anonymous):

no its indeterminate

OpenStudy (freckles):

L is definitely bigger than 1 here

OpenStudy (anonymous):

whats the L? the limit?

OpenStudy (freckles):

I'm calling L=|-infty|=infty since that was the limit we got here L>1 which means it diverges

OpenStudy (anonymous):

but its an infinity over infinity

OpenStudy (freckles):

yes but the top goes to infinity faster than the bottom

OpenStudy (anonymous):

because it has a much larger value right?

OpenStudy (freckles):

x^2>x for large x

OpenStudy (freckles):

\[\frac{x^2}{x}-> \infty \text{ as } x-> \infty \]

OpenStudy (anonymous):

the degree right

OpenStudy (freckles):

yes 2>1

OpenStudy (anonymous):

what if we had \[\sum_{n=1}^{\infty} \frac{ n ^{10} }{ 10^{-n} }\]

OpenStudy (anonymous):

do the n and the 10 cancel out

OpenStudy (freckles):

I would try ratio test again

OpenStudy (anonymous):

like this \[\frac{ n ^{10}}{ 10^{-n} } * \frac{ 10^{-n} }{ n }\]

OpenStudy (freckles):

now you have to replace the n's in the first quotient with (n+1) like we did before

OpenStudy (anonymous):

\[\frac{ n ^{10} }{ 10^{-n}} * \frac{ 10^{-(n+1)} }{ (n+1)^{10} }\]

OpenStudy (freckles):

ok well I was hoping you would replace the n's in the first quotient maybe fine we will find the recipocral of our limit I guess later

OpenStudy (anonymous):

i didnt know you were suppose to replace the n's in both of them

OpenStudy (anonymous):

my mistake

OpenStudy (freckles):

no you aren't

OpenStudy (freckles):

just the first quotient

OpenStudy (anonymous):

let me redo it real quick

OpenStudy (anonymous):

\[\frac{ (n+1)^{10} }{ 10^{-(n+1)}}*\frac{ 10^{-n} }{ n ^{10} }\]

OpenStudy (freckles):

\[L=\lim_{n \rightarrow \infty}|\frac{a_{n+1}}{a_n}|=\lim_{n \rightarrow \infty}|\frac{(n+1)^{10}}{10^{-(n+1)}} \cdot \frac{10^{-n}}{n^{10}}|\] great

OpenStudy (freckles):

you can cancel the 10^(-n) but you would be left with a 10^(-1) on bottom and you also have the ((n+1)/n)^10 to worry about that but that isn't too bad

OpenStudy (anonymous):

so its always the first one that gets the n's replaced by n+1

OpenStudy (freckles):

yes that is what the a_(n+1) divided by a_n tells us to do

OpenStudy (anonymous):

so i'll be left with (n+1)^10 / 10^1 * 1 / n^10

OpenStudy (freckles):

almost you have 10^(-1) on bottom

OpenStudy (freckles):

\[10^{-(n+1)}=10^{-n-1}=10^{-n}10^{-1}\]

OpenStudy (freckles):

the 10^(-n)'s will cancel

OpenStudy (freckles):

you will have 10^(-1) on bottom

OpenStudy (freckles):

\[L=\lim_{n \rightarrow \infty}|\frac{a_{n+1}}{a_n}|=\lim_{n \rightarrow \infty}|\frac{(n+1)^{10}}{10^{-(n+1)}} \cdot \frac{10^{-n}}{n^{10}}| \\= \lim_{n \rightarrow \infty}| \frac{1}{10^{-1}} (\frac{n+1}{n})^{10}| \\ =\lim_{n \rightarrow \infty} 10 (\frac{n+1}{n})^{10}\]

OpenStudy (freckles):

now find the limit as n->infty inside the ( ) part

OpenStudy (freckles):

you basically have n/n since n is really really large n/n=?

OpenStudy (anonymous):

infinity over infinity

OpenStudy (freckles):

do you know what 5/5=? or 5252352/5252352=?

OpenStudy (freckles):

n/n=1

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}10(1)^{10}=10(1)=10\]

OpenStudy (freckles):

L=10>1

OpenStudy (anonymous):

but dont you add the 1+1?

OpenStudy (freckles):

I'm not sure why you are asking that

OpenStudy (freckles):

where does 1+1 come from

OpenStudy (freckles):

pretend n is large n+1 is approximately n you have n/n and n/n=1

OpenStudy (freckles):

this is the same method I used before you basically had 4n^2/n for the previous one and this simplified to 4n

OpenStudy (anonymous):

n+1/n = n/n + 1/n = 1

OpenStudy (freckles):

well that is actually 1+0=1

OpenStudy (freckles):

now 1+1

OpenStudy (freckles):

not*

OpenStudy (anonymous):

okay and since its greater than 1it converges?

OpenStudy (freckles):

no if L>1 then it diverges

OpenStudy (freckles):

you have 10(1)=10>1

OpenStudy (anonymous):

it says converges on the back of the book

OpenStudy (anonymous):

what did we wrong?

OpenStudy (freckles):

\[\sum_{n=1}^{\infty} \frac{n^{10}}{10^{-n}} \text{ diverges }\]

OpenStudy (anonymous):

yeah i think it diverges too i dont know why it says converge

OpenStudy (anonymous):

can we do one more?

OpenStudy (freckles):

did you give the correct problem above

OpenStudy (anonymous):

\[\sum_{n =1}^{\infty} (-1)^{n-1}\frac{ \frac{ 1 }{ 2 }^{n} }{ n^2}\]

OpenStudy (anonymous):

yeah everything is typed in right

OpenStudy (freckles):

ok I'm going to give you a chance to play with it i be right back

OpenStudy (anonymous):

alright for this next one \[-1^{n+1-1}\frac{ \frac{ 1 }{ 2 }^{n+1} }{ (n+10)^2 } * -1^{n-1}\frac{ n^2 }{ \frac{ 1 }{ 2 }^n }\]